# Quantum metrology of low frequency electromagnetic modes with frequency upconverters

### v2

@@ II.3 Quantum amplifier theory
When the RQU is driven by a single microwave tone, it functions as a phase-preserving Caves (1982) electromagnetic amplifier that measures both the low-frequency signal quadratures with equal sensitivity. It is an op-amp mode amplifier Clerk et al. (2010) in the sense that it measures an input state variable (flux) in a lumped-element circuit rather than a traveling wave on a transmission line. The RQU maps the flux variable $\hat{\Phi}$ of the low-frequency circuit onto the microwave output mode $\delta\hat{a}_{out}$.
The upconversion process adds noise as required by the Standard Quantum Limit on amplification, and in this section we show that the RQU can achieve readout at the SQL Clerk et al. (2010). In this readout protocol, the microwave readout tone is resonant ($\Delta=0$), and the low frequency signal appears in $\delta\hat{a}_{out}$ as symmetric sidebands due to the phase modulation of the reflected microwave signal, which is a three-wave mixing process.
In order to evaluate the total noise added in the upconversion process, we calculate fluctuations in the output mode $\delta\hat{a}_{out}$. We use the boundary condition in equation 17 to eliminate the intra-cavity field in equation 24, yielding the equation of motion governing the small quantum fluctuations of the input and output microwave modes, and the low frequency mode:
$\delta\hat{a}_{out}[\omega]=\frac{i\omega-\kappa/2}{i\omega+\kappa/2}\delta\hat{a}_{in}[\omega]+\frac{ig_{0}\bar{a}\sqrt{\kappa}}{(i\omega+\kappa/2)\Phi_{ZPF}}\hat{\Phi}[\omega].$ (26)
@@
where (without loss of generality) we have set the phase of $a_{in}$ so that $\bar{a}$ is real. Equation 27 captures the effects of microwave fluctuations that perturb the state of the low-frequency resonator, including backaction due to fluctuations in the microwave field (proportional to $\delta\hat{a}$ and $\delta\hat{a}^{\dagger}$). Together, equations 26 and 27 describe the two irreducible noise sources in the upconversion process. Since the RQU is functioning as an op-amp (meaning that it measures a state variable rather than a traveling wave), it is more convenient to describe the noise in the upconversion process directly in terms of the voltages and currents in the low-frequency resonator.
We take the limit of low frequencies, where the generated sidebands are within the bandwidth of the microwave resonator, $\omega\ll\kappa/2$. The total output signal at the follow-on microwave amplifier is given by equation 16. In order to recover the flux signal, we noiselessly amplify (with a degenerate JPA) and demodulate using a reference tone at $\omega_{d}$, measuring the microwave phase quadrature $\hat{a}_{out}[\omega]-\hat{a}_{out}^{\dagger}[\omega]$:
$\begin{split}\hat{a}_{out}[\omega]-\hat{a}_{out}^{\dagger}[\omega]=\delta\hat{a}_{in}[\omega]-\delta\hat{a}_{in}^{\dagger}[\omega]+\frac{4ig_{0}\bar{a}}{\sqrt{\kappa}}\frac{\hat{\Phi}[\omega]}{\Phi_{ZPF}}.\end{split}$ (28)
The imprecision fluctuations can be referred back to input currents with equation 28, yielding:
$\hat{I}_{imp}=\frac{\Phi_{ZPF}\sqrt{\kappa}}{4ig_{0}M\bar{a}}\left(\delta\hat{a}_{in}[\omega]-\delta\hat{a}_{in}^{\dagger}[\omega]\right),$ (29)
@@
$\hat{V}_{BA}[\omega]=\frac{8i\omega Mg_{0}Q_{ZPF}|\bar{a}|}{\kappa}(\delta\hat{a}[\omega]+\delta\hat{a}^{\dagger}[\omega]),$ (32)
Equations 32 and 29 show the backaction and imprecision noise contributions arise from the input quadratures $\hat{\delta}_{BA}=\delta\hat{a}_{in}[\omega]+\delta\hat{a}_{in}^{\dagger}[\omega]$ and $\hat{\delta}_{imp}=\delta\hat{a}_{in}[\omega]-\delta\hat{a}_{in}^{\dagger}[\omega]$, respectively. If the input is prepared in a coherent state without additional noise and the output mode is detected without adding noise (e.g. if $\delta\hat{a}_{in},\delta\hat{a}_{in}^{\dagger}$, are sourced from a cold resistor, and a noiseless, phase-sensitive JPA detects $\delta\hat{a}_{out},\delta\hat{a}_{out}^{\dagger}$), the fluctuations of these mode quadratures are described by the (symmetrized) noise spectral densities:
@@
$\bar{S}_{II}=\frac{\Phi_{ZPF}^{2}\kappa}{8|\bar{a}|^{2}g_{0}^{2}M^{2}},$ (36)
$\bar{S}_{VV}=\frac{8|\bar{a}|^{2}g_{0}^{2}M^{2}\omega^{2}Q_{ZPF}^{2}}{\kappa},$ (37)
@@
These spectral densities describe an op-amp mode amplifier operating at the Standard Quantum Limit for a phase-preserving amplifier. The noise impedance of this amplifier is tunable without changing the geometry or input inductance of the device, simply by changing the microwave amplitude $|\bar{a}|$. The total noise added by such an amplifier is the sum of these three contributions, with the op-amp achieving a total added current noise of:
$\bar{S}_{II,tot}=\bar{S}_{II}+\bar{S}_{VV}|\bar{Y}[\omega]|^{2}+2\rm{Re}\left(\bar{S}_{IV}\bar{Y}[\omega]^{*}\right),$ (39)
@@
$R_{\rm noise}=\sqrt{\frac{S_{VV}}{S_{II}}}=\frac{8|\bar{a}|^{2}g_{0}^{2}\omega M^{2}}{\kappa}\frac{Q_{ZPF}}{\Phi_{ZPF}}.$ (41)
@@
$|\bar{a}|^{2}=\frac{R_{b}\kappa}{8g_{0}^{2}M^{2}\omega}\frac{Q_{ZPF}}{\Phi_{ZPF}},$ (42)
@@
This corresponds to an op-amp mode amplifier operating at the SQL.
In reality, we cannot operate the RQU at arbitrarily high microwave drive levels. The upper limit on the value of $|\bar{a}_{in}|^{2}$ will be set by nonlinear terms in the Josephson inductance, leading to Kerr-type nonlinearity, where the effective inductance of the interferometer is a function of amplitude of the currents circulating in the microwave resonator. This introduces terms of the form:
$\hat{H}_{Kerr}=\hbar\Lambda\hat{a}^{\dagger}\hat{a}^{\dagger}\hat{a}\hat{a},$ (44)
where $\Lambda$ quantifies the strength of the quartic nonlinearity in the microwave circuit. Kerr-type nonlinearity has been extensively studied in the context of parametric amplifiers, but the most relevant effect for the RQU is the microwave power-dependent frequency shift. At high microwave powers, this frequency shift causes the microwave resonator to bifurcate. This imposes an upper limit on the microwave power circulating within the RQU Laflamme and Clerk (2011):
$|\bar{a}_{max}|^{2}=\chi|\bar{a}_{bif}|^{2}=\frac{1}{2\sqrt{3}}\frac{\kappa}{\Lambda},$ (45)
where $\chi\lesssim 1$ sets how far below the onset of bifurcation the RQU is operated. This limits the highest noise resistance $R_{max}$ that can be achieved:
$R_{\rm max}=\sqrt{\frac{S_{VV}}{S_{II}}}=\frac{8|\bar{a}_{max}|^{2}g_{0}^{2}\omega M^{2}}{\kappa}\frac{Q_{ZPF}}{\Phi_{ZPF}}.$ (46)
@@ III Measuring untuned input circuits with upconversion
In some applications, especially at low frequencies, resonant input circuits are not practical. Instead, a flux signal is measured in an untuned inductive load such as a magnetometer coil. The sensitivity of this readout is often quantified by its imprecision, uncoupled “energy sensitivity,” which expresses the smallest current signal that can be detected above the imprecision noise for a given inductance Voss (1981). In the case of an untuned circuit, the impedance of the input circuit is so high that backaction noise is insignificant. In this case, the only limit on the noise is set by the fluctuations in the imprecision quadrature $\hat{\delta}_{imp}$, and the final noise temperature is determined by the precision of the phase estimation, as determined by the standard quantum limit on interferometry Pezzé et al. (2007).
@@
where $L_{b}$ is the self inductance of the input coil which couples flux from the low-frequency circuit into the interferometer. The imprecision energy sensitivity can be expressed as a multiple of $\hbar$. The details of the operation of dc SQUIDs limits their imprecision energy sensitivity to ϵgreater-than-or-equivalent-toitalic-$\epsilon\gtrsim\hbar$ (see for example Koch et al. (1981)), but we emphasize that this is not a standard quantum limit. It is nonetheless the appropriate figure of merit for important applications with untuned input circuits.
@@
suggesting that it may outperform the best dc SQUIDs in some untuned applications. To achieve this performance, the RQU must be designed so that this tone power can be applied without approaching junction critical currents or resonator bifurcation Manucharyan et al. (2007).
@@ IV Numerical Estimates of RQU Parameters
In order to inform the design of a practical RQU, we introduce an example low frequency resonator. This example resonator has parameters similar to those that are useful in a variety of precision measurement tasks, e.g. searches for low-mass dark matter candidates Chaudhuri et al. (2018) or high sensitivity measurements of nuclear spin ensembles Sleator et al. (1985). There is substantial flexibility in the design of the RQU which would allow the design to be adapted to higher or lower frequency ranges, or broadband experiments, but the example resonator serves as a benchmark for initial RQU designs.
In order to reach the SQL near the low-frequency resonance, the RQU must be driven with a sufficiently strong microwave tone such that its noise impedance (given by equation 41, with a maximum value set in 46) can match the source impedance of the low frequency resonator, which is entirely determined by its inductance, resonant frequency, and quality factor. We choose a 5 µH inductance as a representative value for a centimeter-scale pickup coil with a few tens of turns (as might be used in a low-mass dark matter experiment searching for electromagnetic Phipps et al. (2020) or nuclear Aybas et al. (2021) interactions). We also note that in principle it is possible to use and $N$-turn step down transformer to lower the pickup coil impedance seen by the RQU by a factor of $1/N$. We do not include this transformer here when evaluating RQU designs, but a transformer might be used to increase the noise impedance further than is achievable with a practical microwave drive tone.
@@ V Two-Tone Microwave Drives
where $a_{\rm drive}$ specifies the amplitude of the two-tone drive, and $\phi_{d}$ sets the phase of the amplitude modulation. Without loss of generality, we can set $\phi_{d}=0$. Solving for the classical amplitude of the field within the microwave resonator yields:
$\bar{a}=a_{\rm max}\cos(\omega_{b}t+\delta)e^{-i\omega_{a}t},$ (53)
where $a_{max}$ captures the rung-up amplitude of the modulated microwave tone:
$a_{\rm max}=a_{\rm drive}\sqrt{\frac{\kappa}{\kappa^{2}+4\omega_{b}^{2}}},$ (54)
@@
We can calculate the equations of motion for $\hat{X}$ and $\hat{Y}$, following the derivation in Clerk et al. (2008). We find that the total noise spectral density of the measured $\hat{X}$ quadrature is given by:
$S_{X}(\omega)=\frac{\gamma/2}{\omega^{2}+(\gamma/2)^{2}}\left(1+2(n_{\rm eq}+n_{\rm bad})\right),$ (58)
where $n_{\rm eq}$ describes the thermal occupation of the low-frequency resonator and $n_{\rm bad}$ describes the spurious backaction on the ideally backaction-free $\hat{X}$ quadrature:
$n_{bad}=\frac{(a_{max}g_{0}\Phi_{ZPF})^{2}}{16\kappa\gamma}\left(\frac{\kappa}{\omega_{b}}\right)^{2}.$ (59)
In the limit of good sideband resolution $\omega_{b}/\kappa\gg 1$, this spurious backaction can be arbitrarily suppressed, corresponding to a backaction-free, or quantum non-demolition (QND) measurement. In this scheme, the RQU can operate at high microwave power levels, reducing imprecision noise on the $\hat{X}$ quadrature, without the backaction penalty present in the single-tone, phase-insensitive mode of operation (as captured by the backaction voltage term in equation 37). This backaction-free measurement mode could allow a quantum speedup in dark matter searches.

### v3

@@ II.3 Quantum amplifier theory
When the RQU is driven by a single microwave tone, it functions as a phase-preserving Caves, (1982) electromagnetic amplifier that measures both the low-frequency signal quadratures with equal sensitivity. It is an op-amp mode amplifier Clerk et al., (2010) in the sense that it measures an input state variable (flux) in a lumped-element circuit rather than a traveling wave on a transmission line. The RQU maps the flux variable $\hat{\Phi}$ of the low-frequency circuit onto the microwave output mode $\delta\hat{a}_{\rm{out}}$.
The upconversion process adds noise as required by the Standard Quantum Limit on amplification, and in this section we show that the RQU can achieve readout at the SQL Clerk et al., (2010). In this readout protocol, the microwave readout tone is resonant ($\Delta=0$), and the low frequency signal appears in $\delta\hat{a}_{\rm{out}}$ as symmetric sidebands due to the phase modulation of the reflected microwave signal, which is a three-wave mixing process.
In order to evaluate the total noise added in the upconversion process, we calculate fluctuations in the output mode $\delta\hat{a}_{\rm{out}}$. We use the boundary condition in equation 17 to eliminate the intra-cavity field in equation 24, yielding the equation of motion governing the small quantum fluctuations of the input and output microwave modes, and the low frequency mode:
$\delta\hat{a}_{\rm{out}}[\omega]=\frac{i\omega-\kappa/2}{i\omega+\kappa/2}\delta\hat{a}_{\rm{in}}[\omega]+\frac{ig_{0}\bar{a}\sqrt{\kappa}}{(i\omega+\kappa/2)\Phi_{\rm{ZPF}}}\hat{\Phi}[\omega].$ (26)
@@
where (without loss of generality) we have set the phase of $a_{\rm{in}}$ so that $\bar{a}$ is real. Equation 27 captures the effects of microwave fluctuations that perturb the state of the low-frequency resonator, including backaction due to fluctuations in the microwave field (proportional to $\delta\hat{a}$ and $\delta\hat{a}^{\dagger}$). Together, equations 26 and 27 describe the two irreducible noise sources in the upconversion process. Since the RQU is functioning as an op-amp (meaning that it measures a state variable rather than a traveling wave), it is more convenient to describe the noise in the upconversion process directly in terms of the voltages and currents in the low-frequency resonator.
We take the limit of low frequencies, where the generated sidebands are within the bandwidth of the microwave resonator, $\omega\ll\kappa/2$. The total output signal at the follow-on microwave amplifier is given by equation 16. In order to recover the flux signal, we noiselessly amplify (with a degenerate JPA) and demodulate using a reference tone at $\omega_{d}$, measuring the microwave phase quadrature $\hat{a}_{\rm{out}}[\omega]-\hat{a}_{\rm{out}}^{\dagger}[\omega]$:
$\begin{split}\hat{a}_{\rm{out}}[\omega]-\hat{a}_{\rm{out}}^{\dagger}[\omega]=\delta\hat{a}_{\rm{in}}[\omega]-\delta\hat{a}_{\rm{in}}^{\dagger}[\omega]+\frac{4ig_{0}\bar{a}}{\sqrt{\kappa}}\frac{\hat{\Phi}[\omega]}{\Phi_{\rm{ZPF}}}.\end{split}$ (28)
The fluctuating terms ($\delta\hat{a}_{\rm{in}}[\omega]-\delta\hat{a}_{\rm{in}}^{\dagger}[\omega]$) in the microwave phase quadrature cause imperfect reconstruction of the low-frequency current at the input of the RQU after microwave amplification and demodulation, which we refer to as imprecision noise. This imprecision noise can be referred back to input currents with equation 28, yielding:
$\hat{I}_{imp}=\frac{\Phi_{\rm{ZPF}}\sqrt{\kappa}}{4ig_{0}M\bar{a}}\left(\delta\hat{a}_{\rm{in}}[\omega]-\delta\hat{a}_{\rm{in}}^{\dagger}[\omega]\right),$ (29)
@@
$\hat{V}_{BA}[\omega]=\frac{8i\omega Mg_{0}Q_{\rm{ZPF}}|\bar{a}|}{\kappa}(\delta\hat{a}[\omega]+\delta\hat{a}^{\dagger}[\omega]),$ (32)
Equations 32 and 29 show the backaction and imprecision noise contributions arise from the input quadratures $\hat{\delta}_{BA}=\delta\hat{a}_{\rm{in}}[\omega]+\delta\hat{a}_{\rm{in}}^{\dagger}[\omega]$ and $\hat{\delta}_{imp}=\delta\hat{a}_{\rm{in}}[\omega]-\delta\hat{a}_{\rm{in}}^{\dagger}[\omega]$, respectively. If the input is prepared in a coherent state without additional noise and the output mode is detected without adding noise (e.g. if $\delta\hat{a}_{\rm{in}},\delta\hat{a}_{\rm{in}}^{\dagger}$, are sourced from a cold resistor, and a noiseless, phase-sensitive JPA detects $\delta\hat{a}_{\rm{out}},\delta\hat{a}_{\rm{out}}^{\dagger}$), the fluctuations of these mode quadratures are described by the (symmetrized) noise spectral densities:
@@
$\bar{S}_{II}=\frac{\Phi_{\rm{ZPF}}^{2}\kappa}{8|\bar{a}|^{2}g_{0}^{2}M^{2}},$ (36)
$\bar{S}_{VV}=\frac{8|\bar{a}|^{2}g_{0}^{2}M^{2}\omega^{2}Q_{\rm{ZPF}}^{2}}{\kappa},$ (37)
@@
These spectral densities describe an op-amp mode amplifier capable of operating at the Standard Quantum Limit for a phase-preserving amplifier. The noise impedance of this amplifier is tunable without changing the geometry or input inductance of the device, simply by changing the microwave amplitude $|\bar{a}|$. The total noise added by such an amplifier is the sum of these three contributions, with the op-amp achieving a total added current noise of:
$\bar{S}_{II,tot}=\bar{S}_{II}+\bar{S}_{VV}|\bar{Y}[\omega]|^{2}+2\mathrm{Re}\left(\bar{S}_{IV}\bar{Y}[\omega]^{*}\right),$ (39)
@@
$R_{\rm noise}=\sqrt{\frac{S_{VV}}{S_{II}}}=\frac{8|\bar{a}|^{2}g_{0}^{2}\omega M^{2}}{\kappa}\frac{Q_{\rm{ZPF}}}{\Phi_{\rm{ZPF}}}.$ (41)
@@
$|\bar{a}|^{2}=\frac{R_{b}\kappa}{8g_{0}^{2}M^{2}\omega}\frac{Q_{\rm{ZPF}}}{\Phi_{\rm{ZPF}}},$ (42)
@@
This corresponds to an op-amp mode amplifier operating at the SQL. Electromagnetic measurement techniques operating at the SQL are available at higher microwave frequencies, but are not presently available for signal frequencies at lower frequencies between dc and VHF. The RQU thus adds valuable capability to the toolbox of techniques available for quantum metrology.
In reality, we cannot operate the RQU at arbitrarily high microwave drive levels. The upper limit on the value of $|\bar{a}_{\rm{in}}|^{2}$ will be set by nonlinear terms in the Josephson inductance, leading to Kerr-type nonlinearity, where the effective inductance of the interferometer is a function of amplitude of the currents circulating in the microwave resonator. This introduces terms of the form:
$\hat{H}_{\rm Kerr}=\hbar\Lambda\hat{a}^{\dagger}\hat{a}^{\dagger}\hat{a}\hat{a},$ (44)
where $\Lambda$ quantifies the strength of the quartic nonlinearity in the microwave circuit. Kerr-type nonlinearity has been extensively studied in the context of parametric amplifiers, but the most relevant effect for the RQU is the microwave power-dependent frequency shift. At high microwave powers, this frequency shift causes the microwave resonator to bifurcate. This imposes an upper limit on the microwave power circulating within the RQU:
$|\bar{a}_{\rm{max}}|^{2}=\chi|\bar{a}_{bif}|^{2}=\frac{1}{2\sqrt{3}}\frac{\kappa}{\Lambda},$ (45)
where $\chi$ is a dimensionless parameter chosen by the operator, which quantifies how far below the onset of bifurcation the RQU is operated, with $\chi\ll 1$ defining the regime of linear microwave response. $|\bar{a}_{bif}|^{2}$ is the average circulating photon number at the onset of bifurcation Laflamme and Clerk, (2011). This limits the highest noise resistance $R_{\rm{max}}$ that can be achieved:
$R_{\rm max}=\sqrt{\frac{S_{VV}}{S_{II}}}=\frac{8|\bar{a}_{\rm{max}}|^{2}g_{0}^{2}\omega M^{2}}{\kappa}\frac{Q_{\rm{ZPF}}}{\Phi_{\rm{ZPF}}}.$ (46)
@@ III Measuring untuned input circuits with upconversion
In some applications, especially at low frequencies, resonant input circuits are not practical. Instead, a flux signal is measured in an untuned inductive load such as a magnetometer coil. The sensitivity of this readout is often quantified by its imprecision, uncoupled “energy sensitivity,” which expresses the smallest current signal that can be detected above the imprecision noise for a given inductance Voss, (1981). In the case of an untuned circuit, the impedance of the input circuit is so high that backaction noise is insignificant. In this case, the only limit on the noise is set by the fluctuations in the imprecision quadrature $\hat{\delta}_{imp}$, and the final noise temperature is determined by the precision of the phase estimation, as determined by the standard quantum limit on interferometry Pezzé et al., (2007).
@@
where $L_{b}$ is the self inductance of the input coil which couples flux from the low-frequency circuit into the interferometer. The imprecision energy sensitivity can be expressed as a multiple of $\hbar$. The details of the operation of dc SQUIDs limits their imprecision energy sensitivity to ϵgreater-than-or-equivalent-toitalic-$\epsilon\gtrsim\hbar$ (see for example Koch et al., (1981)), but we emphasize that this is not a standard quantum limit. It is nonetheless the appropriate figure of merit for important applications with untuned input circuits.
@@
suggesting that it may outperform the best dc SQUIDs in some untuned applications. To achieve this performance, the RQU must be designed so that this tone power can be applied without approaching junction critical currents or resonator bifurcation Manucharyan et al., (2007).
@@ IV Numerical Estimates of RQU Parameters
In order to inform the design of a practical RQU, we introduce an example low frequency resonator. This example resonator has parameters similar to those that are useful in a variety of precision measurement tasks, e.g. searches for low-mass dark matter candidates Chaudhuri et al., (2018) or high sensitivity measurements of nuclear spin ensembles Sleator et al., (1985). There is substantial flexibility in the design of the RQU which would allow the design to be adapted to higher or lower frequency ranges, or broadband experiments, but the example resonator serves as a benchmark for initial RQU designs.
In order to reach the SQL near the low-frequency resonance, the RQU must be driven with a sufficiently strong microwave tone such that its noise impedance (given by equation 41, with a maximum value set in 46) can match the source impedance of the low frequency resonator, which is entirely determined by its inductance, resonant frequency, and quality factor. We choose a 5 µH inductance as a representative value for a centimeter-scale pickup coil with a few tens of turns (as might be used in a low-mass dark matter experiment searching for electromagnetic Phipps et al., (2020) or nuclear Aybas et al., (2021) interactions). We also note that in principle it is possible to use an $N$-turn step down transformer to lower the pickup coil impedance seen by the RQU by a factor of $1/N^{2}$. We do not include this transformer here when evaluating RQU designs, but a transformer might be used to increase the noise impedance further than is achievable with a practical microwave drive tone.
@@ V Two-Tone Microwave Drives
where $a_{\rm drive}$ specifies the amplitude of the two-tone drive, and $\phi_{d}$ sets the phase of the amplitude modulation. Without loss of generality, we can set $\phi_{\rm drive}=0$. Solving for the classical amplitude of the field within the microwave resonator yields:
$\bar{a}=a_{\rm mod}\cos(\omega_{b}t+\delta)e^{-i\omega_{a}t},$ (53)
where $a_{\rm{mod}}$ captures the peak amplitude of the modulated microwave tone:
$a_{\rm mod}=a_{\rm drive}\sqrt{\frac{\kappa}{\kappa^{2}+4\omega_{b}^{2}}},$ (54)
@@
We can calculate the equations of motion for $\hat{X}$ and $\hat{Y}$, following the derivation in Clerk et al., (2008). We find that the $\hat{Y}$ quadrature suffers measurement backaction, with noise spectral density given by:
$\bar{S}_{YY}[\omega]=\frac{\gamma/2}{\omega^{2}+(\gamma/2)^{2}}\left(1+2(n_{\rm eq}+n_{\rm BA}+n_{\rm bad})\right),$ (58)
where $n_{\rm eq}$ describes the thermal occupation of the low-frequency resonator and $n_{\mathrm{BA}}$ is the microwave backaction on the strongly-perturbed $\hat{Y}$ quadrature, and $n_{\rm bad}$ is a spurious backaction term which acts on both quadratures equally. In contrast, the $\hat{X}$ quadrature suffers reduced backaction, with a total noise spectral density (after accounting for imprecision noise added by the microwave readout) of:
$\bar{S}_{XX}[\omega]=\frac{\gamma/2}{\omega^{2}+(\gamma/2)^{2}}\left(1+2(n_{\rm eq}+n_{\rm bad}+n_{\mathrm{imp}})\right),$ (59)
where $n_{\mathrm{imp}}$ represents imprecision noise in the measurement due to noise in the microwave drive and readout. Note that the microwave signal carries no information about the $\hat{Y}$ quadrature, so equation 58 represents only the internal fluctuations of the low-frequency resonator, while equation 59 includes both the internal fluctuations and the additional uncertainty in the $\hat{X}$ quadrature after microwave detection.
In terms of photon numbers, the various noise terms are given by:
$n_{\rm{eq}}=\left(e^{\hbar\omega_{b}/k_{B}T}-1\right)^{-1},$ (60)
$n_{\rm{BA}}=\frac{2(a_{\mathrm{mod}}g_{0}\Phi_{\mathrm{ZPF}})^{2}}{\kappa\gamma},$ (61)
$n_{\rm{bad}}=\frac{(a_{\rm{mod}}g_{0}\Phi_{\rm{ZPF}})^{2}}{16\kappa\gamma}\left(\frac{\kappa}{\omega_{b}}\right)^{2},$ (62)
$n_{\rm{imp}}=\frac{\kappa\gamma S_{0}}{32(a_{\mathrm{mod}}g_{0}\Phi_{\mathrm{ZPF}})^{2}},$ (63)
where $T$ is the equilibrium temperature of the low-frequency resonator, and $S_{0}$ is the noise associated with the microwave readout, with $S_{0}=1$ corresponding to shot-noise limited microwave readout (via, for example, a phase-insensitive JPA operating at the SQL).
The total added noise on the $\hat{X}$ quadrature (not including the equilibrium thermal fluctuations $n_{\mathrm{eq}}$) is:
$n_{\mathrm{add,X}}=n_{\rm{bad}}+n_{\rm{imp}},$ (64)
with the SQL corresponding to $n_{\mathrm{add,X}}=1/2$. The strength of the measurement can be parametrized by $n_{\rm{BA}}$, with the quadrature noises equivalent to:
$n_{\mathrm{add,X}}=\frac{n_{\rm{BA}}}{32}\left(\frac{\kappa}{\omega_{b}}\right)^{2}+\frac{S_{0}}{16n_{\rm{BA}}},$ (65)
In the limit of good sideband resolution ($\omega_{b}/\kappa\gg 1$), this spurious backaction can be arbitrarily suppressed, corresponding to a backaction-free, or quantum non-demolition (QND) measurement. For shot-noise limited readout ($S_{0}=1$), the total noise in the $\hat{X}$ quadrature will evade the SQL as long as the measurement strength reaches $n_{\rm{BA}}\geq 1/8$. Importantly, this backaction evasion scheme can still perform better than the SQL with finite sideband resolution and with non-ideal microwave readout noise. For example, for a very modest sideband resolution ($\omega_{b}/\kappa=2$) and microwave readout noise of a typical HEMT amplifier ($S_{0}\approx 32$), $n_{\mathrm{add,X}}<1/2$ as long as the measurement strength exceeds $n_{\rm{BA}}\approx 4.3$.
In this scheme, the RQU can operate at high microwave power levels, reducing imprecision noise on the $\hat{X}$ quadrature, without the backaction penalty present in the single-tone, phase-insensitive mode of operation (as captured by the backaction voltage term in equation 37). This backaction-free measurement mode could allow a quantum speedup in dark matter searches.